OpenStudy (sbentleyaz):

A 10-ft ladder rests against a vertical wall. If the bottom of the ladder slides away from the wall at a rate of 1 ft/sec, how fast, in ft/sec, is the top of the ladder sliding down the wall, at the instant when the bottom of the ladder is 6 ft from the wall? Answer with 2 decimal places. Type your answer in the space below. If your answer is a number less than 1, place a leading "0" before the decimal point

OpenStudy (phi):

Use Pythagoras to write \[ x^2 + y^2 = 100 \] use "implicit differentiation" to find the derivative of each variable with respect to t (i.e. time): \[ 2 x \frac{dx}{dt} + 2y \frac{dy}{dt} = 0 \] simplify , and substitute in the given numbers: dx/dt = 1 (the sign depends on how we set up the problem. See figure) x=6 \[ y = \sqrt{100-36} = 8\]